# NeuroAMP: A Novel End-to-end General Purpose Deep Neural Amplifier for Personalized Hearing Aids

### v1

@@ 4.3.1 Impact of Model Parameter Scale on Performance
Table II presents the results under the basic (same) parameter setting. With these constraints, the Transformer model consistently exhibited superior performance across most datasets and metrics. It achieved the highest LCC and SRCC values and the lowest MSE, indicating its inherent architectural efficiency and ability to generalize well even with limited parameters.
Table 2: LCC, SRCC, and MSE Results of Different Models on Speech Datasets (basic (same) Parameter Setting)
Metric Model TIMIT TMHINT VCTK (Unseen)
@@
HASPI CNN 0.9707 0.9846 0.0058 0.9863 0.9833 0.0021 0.9314 0.9705 0.0038
LSTM 0.9816 0.9894 0.0037 0.9882 0.9855 0.0018 0.9447 0.9716 0.0031
CRNN 0.9728 0.9788 0.0057 0.9785 0.9736 0.0041 0.9212 0.9437 0.0059
Transformer 0.9814 0.9901 0.0036 0.9903 0.9902 0.0015 0.9548 0.9811 0.0023
HASQI CNN 0.9943 0.9767 0.0013 0.9948 0.9770 0.0014 0.9894 0.9822 0.0020
LSTM 0.9971 0.9869 0.0010 0.9954 0.9784 0.0014 0.9950 0.9924 0.0019
CRNN 0.9902 0.9757 0.0084 0.9878 0.9671 0.0112 0.9851 0.9789 0.0154
Transformer 0.9966 0.9858 0.0007 0.9957 0.9805 0.0010 0.9918 0.9874 0.0011
Table III shows the results after optimizing the parameters. With increased parameter counts, almost all models exhibited consistent performance improvements. Notably, the CNN and LSTM models showed substantial gains in their ability to generalize, particularly on the TMHINT dataset. The CRNN model, which initially struggled under the minimal setting, demonstrated a notable improvement and achieved comparable performance after tuning. Furthermore, the Transformer model achieves the best performance and further improves out-of-domain generalization in the unseen VCTK dataset.
Table 3: LCC, SRCC, and MSE Results of Different Models on Speech Datasets (Optimized Parameter Setting)
Metric Model TIMIT TMHINT VCTK (Unseen)
@@
HASPI CNN 0.9684 0.9822 0.0063 0.9821 0.9801 0.0028 0.9211 0.9619 0.0046
LSTM 0.9822 0.9893 0.0036 0.9872 0.9834 0.0021 0.9435 0.9708 0.0031
CRNN 0.9864 0.9923 0.0026 0.9892 0.9858 0.0016 0.9513 0.9779 0.0025
Transformer 0.9864 0.9927 0.0027 0.9928 0.9923 0.0010 0.9493 0.9810 0.0026
HASQI CNN 0.9941 0.9771 0.0017 0.9929 0.9755 0.0023 0.9870 0.9775 0.0035
LSTM 0.9964 0.9847 0.0018 0.9958 0.9804 0.0023 0.9942 0.9912 0.0026
CRNN 0.9971 0.9864 0.0011 0.9951 0.9794 0.0014 0.9946 0.9920 0.0021
Transformer 0.9979 0.9905 0.0005 0.9965 0.9827 0.0007 0.9941 0.9914 0.0008
@@
Table IV provides a detailed performance evaluation of the CNN, LSTM, CRNN, and Transformer models on the VoiceBank dataset (unseen) under different speech conditions: clean, noisy, and enhanced (using FCN and LSTM). The CRNN model excels in clean and enhanced conditions, achieving high LCC and SRCC scores along with notably low MSE values. The Transformer model also performs well, particularly in clean and enhanced conditions.
@@
Table 4: Detailed evaluation metrics for clean, noisy, and enhanced Speech conditions
Metric Model Clean Enhanced(FCN) Enhanced(LSTM) Noisy
LCC ↑ SRCC ↑ MSE ↓ LCC ↑ SRCC ↑ MSE ↓ LCC ↑ SRCC ↑ MSE ↓ LCC ↑ SRCC ↑ MSE ↓
HASPI CNN 0.9280 0.9654 0.0012 0.8913 0.9209 0.0073 0.8932 0.9199 0.0071 0.9608 0.9781 0.0027
LSTM 0.9457 0.9761 0.001 0.9285 0.9503 0.0045 0.9162 0.9381 0.0053 0.9754 0.9831 0.0016
CRNN 0.9530 0.9819 0.0007 0.9315 0.9543 0.0043 0.9332 0.9543 0.0040 0.9811 0.9883 0.0010
Transformer 0.9427 0.9876 0.0008 0.9465 0.9688 0.0033 0.9145 0.9504 0.0052 0.9781 0.9882 0.0012
HASQI CNN 0.7654 0.3800 0.0088 0.9530 0.9533 0.0026 0.9645 0.9639 0.0006 0.9865 0.9856 0.0020
LSTM 0.9164 0.7983 0.0074 0.9747 0.9791 0.0013 0.9774 0.9795 0.0004 0.994 0.9937 0.0014
CRNN 0.9255 0.8421 0.0055 0.9715 0.9755 0.0011 0.9773 0.9795 0.0004 0.9941 0.9935 0.0012
Transformer 0.9013 0.8062 0.0016 0.9729 0.9693 0.0006 0.9786 0.9781 0.0003 0.9927 0.9922 0.0007
@@
Figure 6: VCTK dataset scatter plots for HASPI scores (a) CNN model architecture (b) LSTM model architecture (c) CRNN model architecture (d) Transformer model architecture
Figure 7: VCTK dataset scatter plots for HASQI scores (a) CNN model architecture (b) LSTM model architecture (c) CRNN model architecture (d) Transformer model architecture
We also utilize scatter plots to visually compare and assess the prediction distribution between the NeuroAMP and the ground truth (NAL-R+WDRC) for various speech datasets. Specifically, we selected two datasets: the TMHINT dataset, which is included in our training data, and the VoiceBank dataset, which is unseen during training. For the evaluation, HASPI and HASQI scores are used as assessment metrics.
Fig. 4 displays scatter plots of the HASPI scores for the TMHINT dataset, separated by each model architecture: (a) CNN, (b) LSTM, (c) CRNN, and (d) Transformer. Fig. 5 shows scatter plots of the HASQI scores for the same dataset. Fig.s 6 and 7 illustrate scatter plots for HASPI and HASQI scores on the VoiceBank dataset, respectively, across all model architectures. Each scatter plot is color-coded to distinguish between different types of data: noisy, enhanced by FCN, enhanced by LSTM, and clean, allowing for a clear visualization of score distributions using HASPI and HASQI scores. The density of points along the diagonal in these plots reflects a high degree of agreement between the model outputs (NeuroAMP) and the ground truth (NAL-R+WDRC), suggesting superior model performance. Specifically, the Transformer and CRNN models exhibit distinctive patterns in these scatter plots. The Transformer model (Fig.s 4d and 7d) shows a particularly dense clustering of points along the diagonal across all types of data, suggesting robust performance and effective handling of variations in audio quality. Moreover, the CRNN model (Fig.s 4c and 7c) shows better alignment with the ground truth but exhibits slightly more dispersion, which may be due to the CRNN’s varying ability to handle certain dataset characteristics or noise levels.
@@ 4.3.2 Performance Evaluation on Music Dataset
Table V presents a comprehensive evaluation of our models using the HAAQI metric to assess music quality against ground-truth scores across two test sets: the Cadenza Challenge and MUSDB18-HQ. We employ LCC, SRCC, and MSE to assess correlation and accuracy. The Cadenza Challenge dataset, which is seen during the training phase, and the unseen MUSDB18-HQ dataset, which is not exposed to the model during training, are used to evaluate the models’ abilities to generalize to new data.
Table 5: Detailed evaluation metrics for Music datasets
@@
CNN 0.9479 0.9305 0.0052 0.9683 0.9674 0.0034
LSTM 0.9660 0.9553 0.0073 0.9840 0.9833 0.0031
CRNN 0.9682 0.9608 0.0051 0.9820 0.9821 0.0024
Transformer 0.9762 0.9738 0.0028 0.9888 0.9892 0.0023
Notably, the performance metrics across these datasets demonstrate that our models, particularly the Transformer and CRNN architectures, can handle unseen scenarios effectively. This is evident from their notable performances in LCC, SRCC, and particularly low MSE values in the MUSDB18-HQ dataset. For example, the Transformer model achieves an LCC of 0.9888 and an SRCC of 0.9892 on the unseen MUSDB18-HQ dataset, indicating notable prediction performance and correlation with the ground truth, despite having no prior exposure during training. These findings confirm the capability of our models to adapt and perform robustly across different types of audio content, such as speech and music, which inherently possess distinct acoustic characteristics. In addition, the models show good generalization performance on the unseen data. Furthermore, the effectiveness of the models across both seen and unseen datasets highlights their potential for practical applications in various real-world scenarios
Next, we utilize scatter plots to visually analyze how well our models predicted HAAQI scores across two different music datasets. In Fig. 9, we show scatter plots for the Clarity-Cadenza Challenge dataset, showing the results for each model architecture: (a) CNN, (b) LSTM, (c) CRNN, and (d) Transformer. Fig. 10 presents scatter plots of HAAQI scores on the MUSDB18-HQ dataset, providing insights into the performance of the same models on this unseen music dataset. The plots indicate that all models performed well, with the Transformer architecture particularly excelling in handling both noisy and clean data, as evidenced by its densely clustered points along the diagonal line. These visualizations align with the numerical results in Table II and demonstrate the models’ ability to maintain high accuracy and consistency across diverse scenarios, including seen and unseen datasets.

### v2

@@ 4.3.1 Impact of Model Parameter Scale on Performance
Table II presents the results under the basic (same) parameter setting, reported as mean ± standard deviation with statistically significant differences indicated. Under these constraints, the Transformer model consistently exhibited superior performance across most datasets and metrics. It achieved the highest LCC and SRCC values and the lowest MSE, indicating its inherent architectural efficiency and ability to generalize well even with limited parameters.
Table 2: LCC, SRCC, and MSE Results of Different Models on Speech Datasets (mean $\pm$ std, basic (same) Parameter Setting).
Metric Model TIMIT TMHINT VoiceBank-DEMAND (Unseen)
@@
HASPI CNN 0.971 $\pm$ 0.006 0.984 $\pm$ 0.003 0.006 $\pm$ 0.001 0.986 $\pm$ 0.003 0.982 $\pm$ 0.004 0.002 $\pm$ 0.001 0.931 $\pm$ 0.018 0.970 $\pm$ 0.002 0.004 $\pm$ 0.001
LSTM 0.982 $\pm$ 0.004 0.989 $\pm$ 0.002 0.004 $\pm$ 0.001 0.988 $\pm$ 0.004 0.984 $\pm$ 0.003 0.002 $\pm$ 0.000 0.944 $\pm$ 0.015 0.971 $\pm$ 0.002 0.003 $\pm$ 0.001
CRNN 0.973 $\pm$ 0.003 0.978 $\pm$ 0.002 0.006 $\pm$ 0.001 0.978 $\pm$ 0.008 0.972 $\pm$ 0.004 0.004 $\pm$ 0.001 0.921 $\pm$ 0.008 0.943 $\pm$ 0.003 0.006 $\pm$ 0.001
Transformer 0.981 $\pm$ 0.006∗‡ 0.990 $\pm$ 0.002∗‡ 0.004 $\pm$ 0.001∗‡ 0.990 $\pm$ 0.002∗†‡ 0.989 $\pm$ 0.003∗†‡ 0.001 $\pm$ 0.000∗†‡ 0.954 $\pm$ 0.019∗†‡ 0.981 $\pm$ 0.002∗†‡ 0.002 $\pm$ 0.001∗†‡
HASQI CNN 0.994 $\pm$ 0.001 0.976 $\pm$ 0.006 0.001 $\pm$ 0.000 0.995 $\pm$ 0.001 0.976 $\pm$ 0.006 0.001 $\pm$ 0.000 0.989 $\pm$ 0.001 0.982 $\pm$ 0.003 0.002 $\pm$ 0.000
LSTM 0.997 $\pm$ 0.000 0.986 $\pm$ 0.002 0.001 $\pm$ 0.000 0.995 $\pm$ 0.001 0.977 $\pm$ 0.007 0.001 $\pm$ 0.000 0.995 $\pm$ 0.000 0.992 $\pm$ 0.001 0.002 $\pm$ 0.000
CRNN 0.990 $\pm$ 0.001 0.975 $\pm$ 0.002 0.008 $\pm$ 0.001 0.988 $\pm$ 0.001 0.965 $\pm$ 0.008 0.011 $\pm$ 0.001 0.985 $\pm$ 0.001 0.978 $\pm$ 0.002 0.015 $\pm$ 0.001
Transformer 0.997 $\pm$ 0.001∗‡ 0.985 $\pm$ 0.003∗‡ 0.001 $\pm$ 0.000∗‡ 0.996 $\pm$ 0.001∗‡ 0.979 $\pm$ 0.006∗‡ 0.001 $\pm$ 0.000∗‡ 0.992 $\pm$ 0.001∗†‡ 0.987 $\pm$ 0.002∗†‡ 0.001 $\pm$ 0.000∗†‡
Significance symbols for the Transformer: vs. CNN ($p<0.05$), vs. LSTM ($p<0.05$), vs. CRNN ($p<0.05$).
Table III shows the results after optimizing the parameters, also reported as mean ± standard deviation and including significance testing. With increased parameter counts, almost all models exhibited consistent performance improvements. Notably, the CNN and LSTM models showed substantial gains in their ability to generalize, particularly on the TMHINT dataset. The CRNN model, which initially struggled under the minimal setting, demonstrated a notable improvement and achieved comparable performance after tuning. Furthermore, the Transformer model achieves the best performance and further improves out-of-domain generalization in the unseen VoiceBank-DEMAND dataset, with several improvements being statistically significant.
Table 3: LCC, SRCC, and MSE Results of Different Models on Speech Datasets (mean $\pm$ std, Optimized Parameter Setting).
Metric Model TIMIT TMHINT VoiceBank-DEMAND (Unseen)
@@
HASPI CNN 0.968 $\pm$ 0.006 0.982 $\pm$ 0.003 0.006 $\pm$ 0.001 0.982 $\pm$ 0.004 0.979 $\pm$ 0.004 0.003 $\pm$ 0.001 0.920 $\pm$ 0.022 0.962 $\pm$ 0.003 0.005 $\pm$ 0.001
LSTM 0.982 $\pm$ 0.003 0.989 $\pm$ 0.002 0.004 $\pm$ 0.001 0.987 $\pm$ 0.005 0.982 $\pm$ 0.003 0.002 $\pm$ 0.000 0.943 $\pm$ 0.013 0.971 $\pm$ 0.002 0.003 $\pm$ 0.001
CRNN 0.986 $\pm$ 0.003 0.992 $\pm$ 0.001 0.003 $\pm$ 0.001 0.989 $\pm$ 0.003 0.984 $\pm$ 0.002 0.002 $\pm$ 0.000 0.951 $\pm$ 0.015 0.978 $\pm$ 0.002 0.003 $\pm$ 0.001
Transformer 0.987 $\pm$ 0.004∗† 0.992 $\pm$ 0.001∗† 0.003 $\pm$ 0.001∗† 0.992 $\pm$ 0.004∗†‡ 0.991 $\pm$ 0.003∗†‡ 0.001 $\pm$ 0.000∗†‡ 0.948 $\pm$ 0.018 0.981 $\pm$ 0.002∗†‡ 0.003 $\pm$ 0.001∗†
HASQI CNN 0.994 $\pm$ 0.001 0.976 $\pm$ 0.006 0.002 $\pm$ 0.000 0.993 $\pm$ 0.001 0.974 $\pm$ 0.005 0.002 $\pm$ 0.000 0.987 $\pm$ 0.001 0.977 $\pm$ 0.002 0.004 $\pm$ 0.000
LSTM 0.996 $\pm$ 0.001 0.984 $\pm$ 0.003 0.002 $\pm$ 0.000 0.996 $\pm$ 0.001 0.979 $\pm$ 0.006 0.002 $\pm$ 0.000 0.994 $\pm$ 0.000 0.991 $\pm$ 0.001 0.003 $\pm$ 0.000
CRNN 0.997 $\pm$ 0.000 0.986 $\pm$ 0.003 0.001 $\pm$ 0.000 0.995 $\pm$ 0.002 0.978 $\pm$ 0.007 0.001 $\pm$ 0.000 0.995 $\pm$ 0.001 0.992 $\pm$ 0.001 0.002 $\pm$ 0.000
Transformer 0.998 $\pm$ 0.000∗†‡ 0.990 $\pm$ 0.001∗†‡ 0.000 $\pm$ 0.000∗†‡ 0.997 $\pm$ 0.002∗†‡ 0.981 $\pm$ 0.005∗‡ 0.001 $\pm$ 0.000∗†‡ 0.994 $\pm$ 0.001 0.991 $\pm$ 0.001 0.001 $\pm$ 0.000∗†‡
Significance symbols for the Transformer: vs. CNN ($p<0.05$), vs. LSTM ($p<0.05$), vs. CRNN ($p<0.05$).
@@
Table IV provides a detailed performance evaluation of CNN, LSTM, CRNN, and Transformer models on the VoiceBank-DEMAND dataset (unseen) under different speech conditions: clean, noisy, and enhanced (using FCN and LSTM). The CRNN model excelled in clean and enhanced conditions, achieving high LCC and SRCC scores along with notably low MSE values. The Transformer model also performs well, particularly in clean and enhanced conditions.
@@
Table 4: Detailed evaluation metrics for clean, noisy, and enhanced speech conditions (mean $\pm$ std)
Metric Model Clean Enhanced Noisy
HASPI CNN 0.932 $\pm$ 0.032 0.963 $\pm$ 0.007 0.001 $\pm$ 0.001 0.893 $\pm$ 0.028 0.917 $\pm$ 0.011 0.007 $\pm$ 0.002 0.961 $\pm$ 0.013 0.977 $\pm$ 0.004 0.003 $\pm$ 0.001
LSTM 0.950 $\pm$ 0.018 0.974 $\pm$ 0.006 0.001 $\pm$ 0.000 0.923 $\pm$ 0.024 0.942 $\pm$ 0.008 0.005 $\pm$ 0.002 0.976 $\pm$ 0.007 0.982 $\pm$ 0.005 0.002 $\pm$ 0.001
CRNN 0.956 $\pm$ 0.025 0.981 $\pm$ 0.004 0.001 $\pm$ 0.000 0.933 $\pm$ 0.019 0.951 $\pm$ 0.006 0.004 $\pm$ 0.002 0.982 $\pm$ 0.003 0.988 $\pm$ 0.002 0.001 $\pm$ 0.000
Transformer 0.951 $\pm$ 0.025 0.986 $\pm$ 0.003 0.001 $\pm$ 0.001 0.931 $\pm$ 0.027 0.958 $\pm$ 0.009∗† 0.004 $\pm$ 0.002 0.979 $\pm$ 0.010 0.987 $\pm$ 0.004 0.001 $\pm$ 0.001
HASQI CNN 0.749 $\pm$ 0.067 0.380 $\pm$ 0.073 0.009 $\pm$ 0.001 0.957 $\pm$ 0.004 0.954 $\pm$ 0.006 0.002 $\pm$ 0.000 0.987 $\pm$ 0.002 0.985 $\pm$ 0.003 0.002 $\pm$ 0.000
LSTM 0.912 $\pm$ 0.023 0.795 $\pm$ 0.043 0.007 $\pm$ 0.001 0.980 $\pm$ 0.002 0.979 $\pm$ 0.002 0.001 $\pm$ 0.000 0.994 $\pm$ 0.001 0.993 $\pm$ 0.001 0.001 $\pm$ 0.000
CRNN 0.922 $\pm$ 0.017 0.838 $\pm$ 0.019 0.006 $\pm$ 0.000 0.979 $\pm$ 0.003 0.979 $\pm$ 0.003 0.001 $\pm$ 0.000 0.994 $\pm$ 0.001 0.993 $\pm$ 0.001 0.001 $\pm$ 0.000
Transformer 0.896 $\pm$ 0.023∗†‡ 0.802 $\pm$ 0.025∗‡ 0.002 $\pm$ 0.000∗†‡ 0.981 $\pm$ 0.003 0.981 $\pm$ 0.003 0.001 $\pm$ 0.000∗†‡ 0.993 $\pm$ 0.002 0.991 $\pm$ 0.002 0.001 $\pm$ 0.000∗†‡
Significance symbols for the Transformer: vs. CNN, vs. LSTM, vs. CRNN $(p<0.05)$. Enhanced results are averaged over both enhancement models.
4.3.2 Hyperparameter Sensitivity Analysis
To further evaluate the stability and generalizability of the proposed method, we conducted two complementary hyperparameter sensitivity analyses—one focusing on model scale (Small, Medium, Large) and one on model depth (1–4 Transformer blocks). For the scale experiment, variants differ in number of layers and hidden units; for the depth experiment, we fixed other parameters and varied only the block count. All models were evaluated on the unseen VoiceBank-DEMAND dataset using the HASPI and HASQI metrics. We report Pearson’s LCC, Spearman’s SRCC, and MSE, each with 95 % confidence intervals.
Figures 4 (scale) and 5 (depth) summarize these results. Across both analyses and all metrics, performance remains remarkably stable, with only marginal gains observed when increasing the model size beyond Medium or adding more than 2–3 blocks. The nearly overlapping confidence intervals indicate that our Transformer-based system is robust to reasonable variations in both size and depth, supporting its practical deployability and generalization capability.
Figure 4: Performance of the Transformer models of varying sizes (Small/Medium/Large) on HASPI and HASQI. (a) LCC, (b) SRCC, and (c) MSE. Solid lines denote the mean; shaded bands indicate $95\%$ confidence intervals.
Figure 5: Performance of the Transformer models with varying number of blocks (1–4) on HASPI and HASQI. (a) LCC, (b) SRCC, and (c) MSE. Solid lines denote the mean; shaded bands indicate 95% confidence intervals.
@@ 4.3.2 Hyperparameter Sensitivity Analysis
Figure 8: VoiceBank-DEMAND dataset scatter plots for HASPI scores (a) CNN model architecture (b) LSTM model architecture (c) CRNN model architecture (d) Transformer model architecture
Figure 9: VoiceBank-DEMAND dataset scatter plots for HASQI scores (a) CNN model architecture (b) LSTM model architecture (c) CRNN model architecture (d) Transformer model architecture
4.3.3 Scatter Plot Analysis
We also utilize scatter plots to visually compare and assess the prediction distribution between the NeuroAMP and the ground truth (NAL-R+WDRC) for various speech datasets. Specifically, we selected two datasets: the TMHINT dataset, used during training, and the VoiceBank-DEMAND dataset, unseen during training. For evaluation, HASPI and HASQI scores are used as assessment metrics.
Fig. 6 displays scatter plots of the HASPI scores for the TMHINT dataset, separated by each model architecture: (a) CNN, (b) LSTM, (c) CRNN, and (d) Transformer. Fig. 7 shows scatter plots of the HASQI scores for the same dataset. Figures 8 and 9 illustrate scatter plots for HASPI and HASQI scores on the VoiceBank-DEMAND dataset, respectively, across all model architectures. In each scatter plot, a red regression line is plotted to indicate the best linear fit, and the shaded area around the line represents the 95% confidence interval, providing a visual measure of the reliability of the linear relationship between the model predictions and the ground truth. Each plot is also color-coded to distinguish between different types of data: noisy, enhanced by FCN, enhanced by LSTM, and clean, allowing for a clear visualization of score distributions using HASPI and HASQI scores. The density of points along the diagonal in these plots reflects a high degree of agreement between the model outputs (NeuroAMP) and the ground truth (NAL-R+WDRC), suggesting superior model performance. Specifically, the Transformer and CRNN models exhibit distinctive patterns in these scatter plots. The Transformer model (Figures 6d and 9d) shows particularly dense clustering of points along the diagonal across all types of data, suggesting robust performance and effective handling of variations in audio quality. Moreover, the CRNN model (Figs 6c and 9c) shows better alignment with the ground truth but exhibits slightly more dispersion, which may be due to the CRNN’s varying ability to handle certain dataset characteristics or noise levels.
@@ 4.3.4 Performance Evaluation on Music Dataset
Table V presents a comprehensive evaluation of our models using the HAAQI metric to assess music quality against ground-truth scores across two test sets: the Cadenza Challenge and MUSDB18-HQ. LCC, SRCC, and MSE are reported, with statistically significant differences indicated. The Cadenza Challenge dataset, which is seen during the training phase, and the unseen MUSDB18-HQ dataset, which is not exposed to the model during training, are used to evaluate the models’ performance and abilities to generalize to new data.
Table 5: Detailed evaluation results for Music datasets (mean $\pm$ std).
@@
CNN 0.95 $\pm$ 0.01 0.93 $\pm$ 0.02 0.005 $\pm$ 0.001 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01 0.003 $\pm$ 0.001
LSTM 0.97 $\pm$ 0.01 0.95 $\pm$ 0.01 0.007 $\pm$ 0.001 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 0.003 $\pm$ 0.001
CRNN 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01 0.005 $\pm$ 0.001 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 0.002 $\pm$ 0.001
Transformer 0.98 $\pm$ 0.01 ∗†‡ 0.97 $\pm$ 0.01 ∗†‡ 0.003 $\pm$ 0.001 ∗†‡ 0.99 $\pm$ 0.00 ∗†‡ 0.99 $\pm$ 0.01 ∗†‡ 0.002 $\pm$ 0.000 ∗†
Significance symbols for Transformer: vs. CNN, vs. LSTM, vs. CRNN $(p<0.05)$.
Notably, the performance metrics across these datasets demonstrate that our models, particularly the Transformer and CRNN architectures, can handle unseen scenarios effectively. This is evident from their notable performances in LCC, SRCC, and particularly low MSE values in the MUSDB18-HQ dataset. For example, the Transformer model achieves an LCC of 0.9888 and an SRCC of 0.9892 on the unseen MUSDB18-HQ dataset, indicating notable prediction performance and correlation with the ground truth, despite having no prior exposure during training. These findings confirm our models’ ability to perform robustly across different audio content types, such as speech and music, which inherently possess distinct acoustic characteristics. In addition, the models show good generalization performance on unseen data. Furthermore, the effectiveness of the models across both seen and unseen datasets highlights their potential for practical applications in various real-world scenarios.
Next, we use scatter plots to visually analyze our models’ predictions of HAAQI scores across two different music datasets. In Fig. 10, we show scatter plots for the Clarity-Cadenza Challenge dataset, showing the results for each model architecture: (a) CNN, (b) LSTM, (c) CRNN, and (d) Transformer. Fig. 11 presents scatter plots of HAAQI scores on the MUSDB18-HQ dataset, providing insights into the performance of the same models on this unseen music dataset. The plots indicate that all models performed well, with the Transformer architecture particularly excelling in handling both noisy and clean data, as evidenced by its densely clustered points along the diagonal line. These visualizations align with the numerical results in Table V and demonstrate the models’ ability to maintain high accuracy and consistency across diverse scenarios, including seen and unseen datasets.